OpenStudy (anonymous):

Is there any particular application For the law of tangents like the law of sines and cosines?

OpenStudy (kc_kennylau):

Are you talking about \(\displaystyle\frac{a-b}{a+b}=\frac{\tan[\frac12(\alpha-\beta)]}{\tan[\frac12(\alpha+\beta)]}\) ?

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

The Wikipedia already has that information: https://en.wikipedia.org/wiki/Law_of_tangents#Application

OpenStudy (kc_kennylau):

"In the time before electronic calculators were available, this method was preferable to an application of the Law of cosines, as this latter law necessitated an additional look-ups in a logarithm table, in order to compute the square root."

ganeshie8 (ganeshie8):

It is equivalent to law of sines, so its just redundant.. eats up extra space in head thats all

OpenStudy (anonymous):

The law of tangents can be used to compute the missing side and angles of a triangle in which two sides a,b and the enclosed angle \gamma are given.

OpenStudy (kc_kennylau):

Congratulations copying from Wikipedia without any quotation mark at all @CO_oLBoY

ganeshie8 (ganeshie8):

\(\dfrac{a}{b} = \dfrac{\sin A}{\sin B} \implies \dfrac{a+b}{a-b} = \dfrac{\sin A + \sin B}{\sin A - \sin B} = \dfrac{\tan((A+B)/2)}{\tan((A-B)/2)}\) law of tangents doesnt provide any additional info that law of sines doest. both are same... its ur choice to memorize whichever is easy for u to..

OpenStudy (anonymous):

So its basically of no use just another form of this sine law

ganeshie8 (ganeshie8):

Exactly^

OpenStudy (anonymous):

@kc_kennylau who said i didnt copied????? btw the same sentence is also written in our text book :P